Find out Ep by the percentage method: Price (SR) Total expenditure (SR) 8 800 10 900
Answer #1

When price is 8, Total expenditure is 800.

Quantity demanded = Total expenditure/Price

Quantity demanded = 800/8 = 100

So,

When price is 8, quantity demanded is 100 units.

When price is 10, Total expenditure is 900.

Quantity demanded = Total expenditure/Price

Quantity demanded = 900/10 = 90

So,

When price is 10, quantity demanded is 90 units.

Calculate the percentage change in quantity demanded -

% change in quantity demanded = [(90 - 100)/100] * 100 = -10%

The percentage change in quantity demanded is -10%

Calculate the percentage change in price -

% change in price = [(10 - 8)/8] * 100 = 25%

The percentage change in price is 25%

Calculate the price elasticity of demand (Ep) -

Ep = % change in quantity demanded/% change in price

Ep = -10/25 = -0.4

Thus,

The Ep by percentage method is -0.4.